OpenStudy (anonymous):

Help please!!!! The specific gravity of iron is 7.87, and the density of water at 4.00C is 1.00g/cm3. You can use this information to find the density of iron. Find the volume occupied by 7.00g of iron

OpenStudy (souvik):

specific gravity of iron=density of iron over density of water so the density of iron=7.87*1=7.87 g/cm^3 now density=mass over volume so the volume =mass over density=7.00/7.87